### v1

Abstract
Cellular traffic forecasting is a critical task that enables network operators to efficiently allocate resources and address anomalies in rapidly evolving environments. The exponential growth of data collected from base stations poses significant challenges to processing and analysis. While machine learning (ML) algorithms have emerged as powerful tools for handling these large datasets and providing accurate predictions, their environmental impact, particularly in terms of energy consumption, is often overlooked in favor of their predictive capabilities.
This study investigates the potential of two bio-inspired models: Spiking Neural Networks (SNNs) and Reservoir Computing through Echo State Networks (ESNs) for cellular traffic forecasting. The evaluation focuses on both their predictive performance and energy efficiency. These models are implemented in both centralized and federated settings to analyze their effectiveness and energy consumption in decentralized systems. Additionally, we compare bio-inspired models with traditional architectures, such as Convolutional Neural Networks (CNNs) and Multi-Layer Perceptrons (MLPs), to provide a comprehensive evaluation. Using data collected from three diverse locations in Barcelona, Spain, we examine the trade-offs between predictive accuracy and energy demands across these approaches.
The results indicate that bio-inspired models, such as SNNs and ESNs, can achieve significant energy savings while maintaining predictive accuracy comparable to traditional architectures. Furthermore, federated implementations were tested to evaluate their energy efficiency in decentralized settings compared to centralized systems, particularly in combination with bio-inspired models. These findings offer valuable insights into the potential of bio-inspired models for sustainable and privacy-preserving cellular traffic forecasting.
keywords:
bio-inspired models, sustainable machine learning, power consumption, spiking neural networks; reservoir computing; federated learning.
1 Introduction
Accurately predicting network traffic and load is essential for improving network management and ensuring efficient operation for telecommunication providers. Reliable forecasting enables proactive resource allocation, effective capacity planning, and improved service quality, especially in high-demand scenarios. The primary challenge lies in maintaining robust connectivity for all devices while avoiding unnecessary resource wastage. Machine learning (ML) algorithms have become indispensable tools for addressing this challenge, offering capabilities such as real-time decision-making, anomaly detection, failure prediction, and dynamic spectrum management. These features make ML a critical component for creating smarter, more sustainable telecommunication networks [1, 2].
This study is motivated by the rapid growth of data generated by telecommunication systems, which is driven by the increasing use of connected devices. The large amounts of data involved require sophisticated predictive models that need significant computational resources to process. As a result, the training of these models often consumes substantial amounts of energy, time, and storage, creating challenges for both efficiency and sustainability.
One major concern is the environmental impact of training such resource-intensive models. As demand for data-driven solutions grows, the energy consumption and carbon emissions from these processes increase. This highlights the need for approaches that can balance strong predictive performance with lower energy consumption.
Figure 1: A architecture for real-time cellular traffic forecasting  [3].
A source of inspiration for this work is the efficiency of the human brain, which processes information using far less energy than modern machine learning systems. For example, training a model like GPT-3 is estimated to consume around 190,000 kWh of energy, whereas the human brain operates on just 12–20 W [4]. By mimicking the brain’s processes, bioinspired models like Spiking Neural Networks (SNNs) offer a promising way to make machine learning more energy-efficient.
In this study, we explore the use of bioinspired SNNs and Reservoir Computing (RC) with Echo State Networks (ESNs) for cellular traffic forecasting. We examine several types of Leaky Integrate-and-Fire (LIF) neurons in SNNs, including Lapicque, Leaky, RLeaky, Synaptic, and Alpha neurons [4], as well as ESNs [5]. These models are compared with traditional machine learning methods, such as Multi-Layer Perceptrons (MLPs) and Convolutional Neural Networks (CNNs). Both centralized and federated learning settings are used in the evaluation to account for modern distributed computation approaches. Federated learning, in particular, reduces the need for centralized data storage, offering a more energy-efficient way to train models.
This study evaluates the trade-offs between predictive accuracy and energy efficiency. While energy-efficient models often show reduced prediction performance, this work aims to explore ways to strike a balance between sustainability and accuracy. To achieve this, we use a Sustainability Index [6] that combines key metrics like power consumption, prediction error, and dataset size to provide a comprehensive evaluation of model efficiency.
To the best of our knowledge, this is the first study to compare the energy consumption of bio-inspired and traditional models for cellular traffic forecasting, evaluated across both centralized and federated learning settings.
We further explore the power efficiency of federated learning, which was initially introduced to enhance privacy in machine learning. Recent studies [7] indicate that federated learning can also reduce power consumption, offering a dual benefit of improved privacy and energy efficiency. This makes it a promising framework for sustainable and privacy-preserving machine learning.
The results aim to guide researchers and practitioners toward developing machine learning models that balance high performance with lower energy use. The rest of the paper is structured as follows: Section 2 re/views related work on Spiking Neural Networks (SNNs), Reservoir Computing (RC), and Federated Learning. Section 3 describes the dataset and research problem, along with the models and evaluation methods. Section 4 explains the experimental setup and methodology. Section 5 presents the results and analysis. Section 6 discusses the findings, challenges, and future directions. Finally, Section 7 concludes by summarizing the key outcomes and contributions of the study.
2 Related Work
2.1 Traffic Prediction
Recent advancements in cellular traffic forecasting have demonstrated the effectiveness of machine learning and federated learning approaches in managing 5G base stations. Perifanis et al. (2023) [8] introduced a federated learning framework for traffic prediction in 5G base stations, highlighting the model’s ability to preserve data privacy while enabling distributed training across multiple stations. Their work emphasized the importance of efficient time-series forecasting to optimize resource allocation and reduce energy consumption in large-scale cellular networks.
Building on this foundation, Pavlidis et al. (2024) [9] extended the concept by applying federated models to predict traffic across a larger number of base stations, while also investigating the impact of fine tuning several experimental parameters. They demonstrated how federated learning could improve forecasting accuracy and reduce unnecessary communication overheads.
Zhao et al. (2023) proposed the FL-DeepAR framework, which uses deep autoregressive networks for federated traffic prediction. This approach addressed statistical heterogeneity and incorporated advanced gradient-based aggregation techniques, achieving superior prediction performance on real-world datasets [10].
Similarly, Guesmi et al. (2024)  [11] proposed advanced predictive modeling techniques to enhance traffic forecasting in emerging cellular networks. Their research leveraged a combination of deep learning and federated learning algorithms to address the challenges of time-series forecasting in dynamic and heterogeneous environments, further reinforcing the potential of distributed machine learning in telecommunications.
Lee et al. (2024)  [12] explored federated learning applications beyond traffic forecasting, focusing on comprehensive mobile network management. Their study emphasized the integration of federated models into resource allocation and small-cell base station management, showcasing the broader impact of this technology on network efficiency.
Gao et al. (2024) developed FedCE, a communication-efficient federated learning framework that reduced communication overhead while improving prediction accuracy using gradient compression and adaptive aggregation strategies. The results highlighted significant improvements in prediction accuracy with optimized resource utilization [13].
Additionally, Lee et al. (2024) proposed a personalized federated learning approach for mobile traffic prediction, which utilized layer-wise aggregation and adaptive layer freezing to improve prediction accuracy and communication efficiency. This method demonstrated its potential for reducing overhead and ensuring scalability in dynamic network environments [14].
Lastly, Kalam et al. (2023) introduced a federated learning approach for 5G traffic forecasting, demonstrating significant improvements in accuracy over traditional centralized methods while ensuring privacy preservation across distributed networks [15].
Collectively, these studies underscore the transformative potential of federated learning and complementary machine learning approaches in cellular traffic forecasting. They offer scalable, privacy-preserving, and resource-efficient solutions to optimize the performance and sustainability of 5G and beyond networks, paving the way for the next generation of mobile communication systems.
2.2 SNN in Traffic Prediction
Recent advancements in spiking neural networks (SNNs) demonstrate their potential for efficient and accurate network traffic prediction, particularly in energy-constrained environments. In our previous work  [3] we explored SNNs in the context of cellular traffic forecasting, emphasizing their ability to reduce power consumption compared to conventional machine learning models. By leveraging the bio-inspired architecture of SNNs, the study addressed the growing demand for environmentally sustainable machine learning solutions in telecommunications. We highlighted how SNNs effectively capture temporal dynamics in network traffic patterns, enabling precise predictions with significantly reduced energy footprints.
Similarly, Kang et al. (2024)  [16] investigated the use of SNNs in neuromorphic systems for traffic hotspot prediction in network-on-chip (NoC) architectures. Their study focused on utilizing liquid state machines (LSMs)—a specific form of reservoir computing based on spiking dynamics—to predict and mitigate traffic congestion hotspots in NoC environments. The findings showcased how SNN-based approaches not only improve computational efficiency but also enhance routing strategies, contributing to a more energy-efficient processing pipeline.
Together, these studies underscore the transformative potential of SNNs in network traffic prediction, providing scalable, low-power solutions for real-time applications while maintaining high predictive accuracy.
2.3 RC in Traffic Prediction
Reservoir computing has emerged as a promising framework for cellular traffic forecasting due to its ability to process complex temporal data efficiently while minimizing computational power consumption. Li et al. (2021)  [17] introduced a deep multi-reservoir regression learning network tailored for cellular traffic prediction in multi-access edge computing environments. This innovative approach leverages the strengths of reservoir computing to provide real-time and accurate predictions, demonstrating its scalability and efficiency in handling dynamic traffic patterns while maintaining low energy overheads.
Further advancing the field, Zhang and Vargas (2023)  [18] applied reservoir computing in the context of 5G and beyond networks. Their study emphasized how reservoir computing architectures enhance forecasting accuracy while addressing sustainability challenges in mobile networks. By leveraging innovative learning techniques, their work provides actionable insights for integrating reservoir computing frameworks into modern telecommunications infrastructures, ensuring both predictive performance and energy efficiency.
Collectively, these studies underline reservoir computing’s transformative potential for cellular traffic forecasting, particularly in energy-constrained and dynamic environments like 5G networks. Its ability to balance computational efficiency with prediction accuracy makes it a critical tool for future network management and optimization.
2.4 Federated Bio Inspired in Traffic Prediction
Federated Learning (FL) addresses privacy concerns in mobile networks but faces challenges like high energy consumption, communication overhead, and convergence issues. Aouedi et al. (2023) [19] introduced HFedSNN, a Hierarchical Federated Learning (HFL) model using Spiking Neural Networks (SNNs) to overcome these issues. HFedSNN reduces energy consumption by 4.3x, communication overhead by 26x, and improves accuracy by 4.48% compared to FL with ANNs (FedANN), making it highly efficient for non-IID data scenarios.
Similarly, Bacciu et al. (2021) [20] introduced Federated Reservoir Computing (FRC) that leverages bio-inspired models like Echo State Networks (ESNs) for low-power, distributed temporal data processing. ESNs excel in handling time-series data while minimizing energy use, offering scalable, privacy-preserving solutions.
Together, HFL with SNNs and FRC provide a powerful framework for robust, energy-efficient, and scalable AI in edge computing environments.
Our work advances the field of cellular traffic prediction by addressing critical challenges in energy efficiency and sustainability, particularly in the context of bioinspired models. While prior research has demonstrated the potential of federated learning to improve prediction accuracy and preserve privacy, much of the focus has been on traditional machine learning methods. Similarly, bioinspired models, such as SNNs and RC, have been explored for their energy efficiency in specific use cases, but their integration with federated learning frameworks remains underexplored. Existing approaches often overlook the trade-offs between energy consumption, predictive performance, and environmental impact.
aOur study distinguishes itself by combining SNNs and RC models within federated learning environments to assess their predictive capacity and energy efficiency. Unlike prior studies that prioritize either centralized configurations or accuracy alone, we employ a Sustainability Index (See Section 3) to evaluate the trade-offs across multiple dimensions, including power consumption, prediction error, and data size. This comprehensive approach, coupled with comparative evaluations against traditional architectures such as MLPs and CNNs, provides deeper insights into how bioinspired models can support sustainable and scalable solutions for cellular traffic forecasting. By leveraging real-world datasets and consistent experimental setups, our work offers a holistic framework for advancing energy-efficient and privacy-preserving network management practices.
3 System Model
This section outlines the framework used to evaluate the performance and energy efficiency of the different models and settings. This includes details about the dataset, the problem formulation, and the model architectures implemented in both centralized and federated learning setups.
3.1 Dataset
In our analysis, we utilized a real-world dataset with LTE traffic measurements that was gathered from three different areas in Barcelona, Spain. The dataset was collected under SUPERCOM initiative, [^1] https://supercom.cttc.es/ from Downlink Control Information (DCI) messages, which are transmitted through the PDCCH every Transmission Time Interval (TTI). Then, the raw data are aggregated into two minutes intervals. Each measurement contains eleven features regarding network traffic (see Table  3.1) and the downstream task is to predict the five of them, i.e. RNTI Count, UpLink, DownLink, RB Up and RB Down, using historical measurements for the next timestep.
[…]
Specifically, the measurements are gathered in the following locations:
  • LesCorts: A residential area near Camp Nou Football Stadium. Measurements at this location comprise 12 days (2019-01-12 to 2019-01-24).
  • PobleSec: A residential and touristic area near the historic center, the exhibition centre and the port. Measurements at this location comprise 28 days (2018-02-05 to 2018-03-05).
  • ElBorn: A touristic area in the downtown of the city. It is characterized by having a lot of shops and nightlife. Measurements at this location comprise 7 days (2018-03-28 to 2018-04-04).
As showed in [8], the dataset presents non Independent and Identically Distributed (non-IID) distibution, with the presence of three different skews, i.e. quantity, quality and temporality. This differentiation among respective BS is depicted in Fig. 3.1, indicating a challenging ML task.
{figurehere}[Uncaptioned image]
Loss Convergence of SNN models  [3]
3.2 Problem Statement
We formulate the problem in the context of time-series forecasting. Consider a dataset $\mathbf{X}\in\mathbb{R}^{N\times d}$, where $N$ is the number of observations in the dataset, and $d$ is the number of measurements recorded at each time step. In our context, these measurements include traffic-related metrics such as the uplink and downlink traffic and the number of connected devices at each time step. Given a window of $S$ past observations $\mathbf{X}_{t-S+1:t}=\{\mathbf{x}_{t-S+1},\mathbf{x}_{t-S+2},\ldots,\mathbf{x}_{t}\}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ is the measurement vector at time $i$, the goal is to predict the measurement vector $\mathbf{x}_{t+1}\in\mathbb{R}^{d^{\prime}}$ at the next time step $t+1$, where $d^{\prime}\leq d$ represents the dimensions of interest for forecasting.
The objective can be formalized as learning a function $f$ that maps a sequence of $S$ past observations to a future value:
$\hat{\mathbf{x}}_{t+1}=f(\mathbf{X}_{t-S+1:t};\theta),$ (1)
where $\hat{\mathbf{x}}_{t+1}\in\mathbb{R}^{d^{\prime}}$ is the predicted measurement vector for the next time step, and $\theta$ represents the parameters of the forecasting model. The model aims to minimize the forecasting error between the predicted measurements $\hat{\mathbf{x}}_{t+1}$ and the actual measurements $\mathbf{x}_{t+1}$.
3.3 ML Models
We evaluate the performance of several model architectures within our framework. Specifically, we first employ two traditional baseline learning models, i.e., MLP and CNN as benchmarks. To facilitate a comprehensive comparative analysis, we then selected several distinct Leaky-and-Integrate Fire (LIF) spiking neuron types: the Lapicque’s original Neuron, as well as its streamlined counterpart, the Leaky Model, alongside the Recurrent Leaky Neuron, the Synaptic Conductance-Based LIF Neuron Model, and the Alpha Neuron Model and a Reservoir Computing (RC), Echo State model. Specifically, we utilize the following model architectures:
  • Multi-Layer Perceptron (MLP): A MLP belongs to the class of feed-forward artificial neural networks (ANNs). MLP can not operate directly on matrix representations, thus requiring time-series to be flattened to one-dimensional vectors. In our case, we transform the $(S,d)$ two-dimensional array used as input into an one-dimensional vector to $(S\times d)$. In our architecture, we utilize a single hidden layer with 128 neurons.
  • Convolutional Neural Network (CNN): CNNs are a type of deep neural networks, highly effective in analyzing data with a grid-like topology. In multivariate time-series, convolution operations can effectively utilize structural covariance. The implemented CNN takes a two-dimensional matrix of size $(S,d)$ as input and feeds it to four two-dimensional convolutional layers. The output is then propagated to a two-dimensional average pooling layer and finally to a fully connected layer. For the final layer, similar to MLP, we utilize a single hidden layer with 128 units.
  • Lapicque Neuron (Original LIF Neuron): The Lapicque neuron implements the fundamental circuit approach, originating from biology. Ion channels of a physical neuron’s membrane are modeled with a resistor and the membrane itself is modeled with a capacitor. We check the potential ($U$) in the capacitor, and if this potential reaches a predefined threshold, the neuron fires an output spike. We also choose a mechanism that is activated when a spike is fired to reset our neuron. Thus, the tunable parameters in this approach are the resistance ($R$), capacitance ($C$), threshold, the number of spikes for each input ($I_{in}$) and the reset mechanism. Formally, Lapicque neuron [4, 21] is modeled through the following function:
    $U[t+1]=(1-\frac{1}{RC})\cdot U[t]+\frac{1}{RC}\cdot I_{in}[t+1]\cdot R,$ (2)
    When the resulting value of $U$ is higher than the specified threshold, the neuron will fire a spike.
  • Leaky Neuron: The initial description of the LIF model, as described in [4] and [21], is distinguished by a broad set of hyperparameters. As the model is scaled to support a universal SNN architecture, managing these parameters becomes challenging. For this type of neuron, we use the following equation:
    $U[t+1]=\beta\cdot U[t]+(1-\beta)\cdot I_{in}[t+1],$ (3)
    where $R=1$ and $\beta=(1-\frac{1}{C})$, represeting a decay rate of the membrane potential. The spikes in the input increase the membrane potential, and when this value exceeds the threshold, a spike is fired in the output.
  • Recurrent Leaky (RLeaky) Neuron: This implementation, as discussed in [4, 22], employs a logic similar to that of Leaky neuron, yet it uniquely scales the output spikes of a neuron by a factor of $V$ and reincorporates them into the input. Specifically, spikes produced by a layer of spiking neurons are forwarded to the following fully connected (or convolutional) layer, where they are combined with the incoming input before being processed by the spiking neuron model.
  • Synaptic Based Neuron: In the neuron models discussed previously, the premise has been that a voltage spike upon input instantaneously elevates the synaptic current, subsequently influencing the membrane potential. However, in practical scenarios, a spike initiates a progressive discharge of neurotransmitters from the pre-synaptic to the post-synaptic neuron. The synaptic conductance-based [23, 21] LIF model accommodates the slow temporal progression of the input current. Thus, the equation is transformed to:
    $U[t+1]=\beta\cdot U[t]+I_{syn}[t+1],$ (4)
    where $I_{syn}[t+1]=\alpha\cdot I_{syn}[t]+I_{in}[t+1]$, $\alpha$ is the decay rate of the synaptic current, and $\beta$ is the decay rate of the membrane potential.
  • Alpha Neuron: This model, as conceptualized in [24], resembles a type of filter. Upon receiving an input spike, it is convolved with the filter, eliciting a change in the membrane potential. The configuration of the filter can be as simple as an exponential curve, or involve multiple exponentials. The attractiveness of these models lies in their ability to easily integrate a range of characteristics, such as refractoriness and threshold adjustment, by embedding these features within the design of the filter itself.
  • Reservoir Computing Echo State: Reservoir computing leverages a fixed, randomly initialized dynamic reservoir to project input data into a high-dimensional feature space. The Echo State Network (ESN) is a prominent implementation of this paradigm, particularly suited for sequential data processing. The reservoir consists of recurrently connected nodes, which transform the input time-series into a rich set of temporal features. Unlike traditional neural networks, the reservoir weights remain fixed, and only the output layer is trained, significantly reducing the computational complexity. In our implementation, the input data $(S,d)$ is mapped to the reservoir with a dimensionality of $N_{r}$ nodes. The output is then computed as a linear combination of the reservoir states and passed to a fully connected layer for final predictions. This architecture facilitates efficient processing of multivariate time-series with minimal parameter tuning and low energy consumption.
3.4 Evaluation Metrics
The primary metric used to quantify prediction error for the time-series forecasting task is the normalized root mean squared error ($\mathrm{NRMSE}$), defined as:
$\mathrm{NRMSE^{(k)}}=\frac{1}{\overline{x}}\sqrt{\frac{\sum_{t=1}^{T}\left(\sum_{i=1}^{d^{\prime}}\left(\hat{x}_{t,i}-x_{t,i}\right)^{2}\right)}{Td^{\prime}}},$ (5)
where $\overline{x}$ represents the mean of the target observations, computed as:
$\overline{x}=\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\left(\sum_{i=1}^{d^{\prime}}x_{t,i}\right),$ (6)
and $\hat{x}_{t,i}$ and $x_{t,i}$ denote the predicted and actual values at time $t$ for feature $i$, $T$ is the total number of time steps, and $d^{\prime}$ is the number of features.
In addition to $\mathrm{NRMSE}$, we also employ other common error metrics to evaluate model performance:
  • Mean Absolute Error (MAE):
    $\mathrm{MAE^{(k)}}=\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\sum_{i=1}^{d^{\prime}}\left|\hat{x}_{t,i}-x_{t,i}\right|,$ (7)
  • Root Mean Square Error (RMSE):
    $\mathrm{RMSE^{(k)}}=\sqrt{\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\sum_{i=1}^{d^{\prime}}\left(\hat{x}_{t,i}-x_{t,i}\right)^{2}},$ (8)
  • Mean Square Error (MSE):
    $\mathrm{MSE^{(k)}}=\frac{1}{Td^{\prime}}\sum_{t=1}^{T}\sum_{i=1}^{d^{\prime}}\left(\hat{x}_{t,i}-x_{t,i}\right)^{2}.$ (9)
Beside the accuracy of the implemented algorithms, we also utilize the Sustainability Indicator $\mathcal{S}$ introduced in [6] to assess the resulting trade-off between the prediction error and the energy consumption. $\mathcal{S}$ enables a fair comparison among different model architectures and can be calculates using the following formula:
$\mathcal{S}=(1+E_{\text{Val}})^{a}\times(1+C_{\text{Tr}})^{b}\times(1+D)^{c},$ (10)
where $E_{\text{Val}}$ is the validation error (in this work, we consider NMRSE), $C_{\text{Tr}}$ represents the total energy consumed for model training in Watt Hours (Wh) and $D$ is the data size to be transmitted in kB.
4 Experimental Setup
To ensure the integrity and validity of our comparative analysis, we maintained consistency across all experimental settings. The architecture of the SNN models was standardized to include a single hidden layer with 128 neurons and five output neurons corresponding to the target variables.
In our SNN models, both the input and output layers utilize the Subtract-Reset Mechanism. This means that when a neuron’s membrane potential exceeds the threshold, a spike is emitted, and the membrane potential is reduced by the threshold value ($V_{\text{mem}}=V_{\text{mem}}-\text{threshold}$).
In the output layer, however, neurons lack a reset mechanism—a deliberate design choice to better address the requirements of the time-series forecasting task. This configuration ensures that the predicted output corresponds directly to the membrane potential, avoiding distortions caused by resets.
Additionally, input spikes are encoded directly from the input data at specific timesteps, enabling the network to effectively convert non-spiking data into a format compatible with spiking mechanisms. This approach enhances the model’s ability to process time-series data within the spiking neural network framework.
The size of the dataset used for training the model is also a crucial factor. Training on a smaller dataset inevitably results in lower power consumption compared to a larger one. This relationship is captured by the Sustainability Index (S), which includes dataset size as a parameter. To ensure consistency, we used the same dataset and size across all experiments.
For implementation, we utilized the PyTorch library [25] for traditional models such as MLPs and CNNs. The SNN models were implemented using the snnTorch library [4], which is tailored for gradient-based learning and provides pre-designed spiking neuron models that seamlessly integrate as recurrent activation units. Energy consumption during the training process was measured using the CarbonTracker Python module [26]. CatbonTracker is a tool designed to estimate and monitor the carbon emissions associated with computational processes, helping organizations and researchers assess and reduce their environmental impact.
Centralized training was conducted over 150 epochs with a batch size of 128. In the federated learning experiments, we used 10 federated steps with 3 local epochs per client with a central server, simulating the setup to maintain consistency.
The experiments were conducted on a Windows 11 workstation equipped with an AMD Ryzen 7 6800HS CPU and 16 GB of memory.
5 Experimental Evaluation
5.1 Predictive Performance
5.1.1 Modeling of Input Data
Selecting the correct number of spiking timesteps is crucial for effectively encoding input data in SNN models. The number of timesteps determines how the input data is represented as spikes, influencing the network’s ability to process and learn from the encoded information. Using too few timesteps may result in a loss of critical details during the encoding process, leading to poor model performance. On the other hand, an excessive number of timesteps can overcomplicate the encoding, increasing computational costs and energy consumption without improving the predictive capacity of the model. This challenge highlights the importance of fine-tuning the timestep configuration to achieve an efficient and accurate representation of the input data.
In our experiments, we investigated how varying the number of timesteps used to encode input data impacts the predictive performance of the models. Specifically, we tested configurations with 1, 7, 10, 50, 70, and 100 timesteps to identify the optimal setup for maximizing accuracy.
As shown in Figure 5.1.1, the results indicate significant variations in performance depending on the chosen timestep length. Optimal performance was achieved at 10 timesteps for some models and 7 timesteps for others. Beyond these values, increasing the number of timesteps did not improve accuracy and, in some cases, resulted in performance degradation. These findings highlight the importance of selecting an appropriate timestep configuration to balance information completeness, computational efficiency, and predictive accuracy.
{figurehere}[Uncaptioned image]
Test Loss for different number of spikes per input
Based on these observations, we selected 7 and 10 timesteps for our examples, as they demonstrated the best performance. This choice minimizes the increase in computational complexity, keeping the calculations as simple as possible to reduce power consumption, while maintaining predictive performance at an acceptable level.
5.1.2 Centralized Learning
The evaluation of model performance and sustainability across centralized configurations reveals critical insights into the trade-offs between predictive accuracy, energy efficiency, and overall sustainability. Figures 2 and 3 present the performance of various models using metrics such as NRMSE, MSE, RMSE, and MAE. The results show that model performance varies significantly depending on the selected metric. To ensure consistency and comparability, NRMSE was chosen as the primary metric, as it normalizes errors relative to the data scale, providing a more realistic and interpretable estimation.
From the data in Tables 5.1.2 and 5.1.3, as well as the referenced figures, we analyzed the performance of the models for 7 and 10 timesteps. The findings reveal that the Leaky architecture is the only model that performs better with 7 timesteps. In contrast, all other models achieve superior performance and lower error rates with 10 timesteps, highlighting the importance of selecting an optimal timestep configuration for each model to balance computational efficiency and predictive accuracy.
At first glance, the Sustainability Index used to evaluate the models does not reveal a straightforward trade-off. Reducing model complexity often leads to lower power consumption but can result in higher error rates, which negatively impacts the Sustainability Index (where a lower value of S indicates greater sustainability). Ultimately, the Sustainability Index reflects the balance between energy efficiency and performance, and the critical question becomes how much power consumption can be reduced without a significant sacrifice in predictive accuracy.
Among the centralized configurations, the Leaky Neuron stands out as the most sustainable model during the training process. Despite its higher prediction error compared to other models, its remarkably low power consumption compensates for this limitation, resulting in the lowest S. This makes the Leaky Neuron the best overall performer in terms of energy efficiency and environmental impact.
In contrast, the Alpha Neuron demonstrates the best predictive accuracy with 10 timesteps, indicating its ability to effectively handle the task. However, this performance comes at the expense of significantly higher power consumption, as the model operates as a computationally intensive filter. Consequently, it has the worst S among all evaluated models.
The Lapicque Neuron exhibits the poorest performance, with the highest RMSE and a power consumption higher than both the MLP and CNN models. This combination of low accuracy and high energy demand results in the lowest S, making the Lapicque Neuron the least favorable choice for centralized configurations.
The RLeaky model presents a particularly intriguing case. While it achieved its best performance at 10 timesteps with improved predictions and a lower NRMSE, this came at the cost of significantly higher power consumption during training. This trade-off highlights the challenge of balancing accuracy and energy efficiency, making the RLeaky model less suitable for applications where sustainability is a critical consideration.
The Echo State model demonstrates a better balance between accuracy and power consumption. Ranking second in terms of sustainability, it outperforms models such as the Lapicque Neuron, MLP, and CNN. Its architecture leverages energy efficiency without substantial performance compromises, making it a promising alternative for resource-constrained applications.
Finally, the Synaptic model, which aims to replicate the complex functions of brain cells, demonstrates significant drawbacks in terms of resource requirements. It consumes substantially more power and produces higher NRMSE compared to the MLP and CNN. This result suggests that the additional biological realism in the Synaptic model comes at the expense of computational efficiency and predictive accuracy, raising concerns about its practicality in real-world applications.
Collectively, these findings emphasize the importance of carefully balancing energy consumption, predictive performance, and sustainability when selecting models for specific tasks. While models like the Leaky Neuron offer an optimal trade-off for energy efficiency and environmental impact, others, such as the Alpha and RLeaky neurons, highlight the inherent trade-offs between accuracy and energy consumption. These insights provide valuable guidance for model selection in diverse application contexts.
Figure 2: Energy vs Prediction 7 Timesteps
Figure 3: Energy vs Prediction 10 Timesteps
[…]
Centralized Cumulative results on NRMSE, Energy Consumption and Sustainability of considered models. Model NRMSE Consumption (Wh) $\mathcal{S}$ MLP 1.3465 0.44 25.401 CNN 1.3272 1.403 29.993 Echo State (RC) 1.42877 0.431 25.637 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input Leaky 1.3168 1.3531 0.0742 0.0785 22.960 23.109 Lapicque 1.9429 1.7641 1.4044 1.797 32.414 33.377 RLeaky 1.3699 1.3055 1.4343 1.875 30.302 31.724 Synaptic 1.3896 1.42 1.3318 2.33 29.956 33.838 Alpha 1.2832 1.2751 2.0561 4.343 32.265 34.075
5.1.3 Federated Learning
The graphs in Figure 4 provide a comprehensive comparison of energy consumption, predictive performance, and sustainability of federated learning configurations across various machine learning models. In those graphs the number in the of the model name means the number of timesteps, e.g. the Leaky7 means the Leaky model with 7 timesteps.
The first graph illustrates the energy consumption of the models under the two paradigms, highlighting the efficiency advantage of centralized learning, which generally requires less energy due to localized computation. In contrast, federated learning models exhibit higher energy consumption, largely due to the communication overhead inherent in decentralized training.
The second graph compares the Normalized Root Mean Square Error (NRMSE) of the models, offering insight into their predictive accuracy. Centralized models typically achieve better performance, reflected in their lower NRMSE values, as they can directly access and process all data. Federated models, while exhibiting higher error rates, still maintain reasonable predictive performance, making them suitable for scenarios where privacy and decentralization are prioritized.
The third graph examines the trade-off between energy consumption and NRMSE, visualizing the balance between resource efficiency and predictive accuracy. This trade-off is critical for assessing the suitability of different learning paradigms in energy-constrained environments, such as edge devices or mobile applications. Federated learning’s higher energy costs and slightly reduced accuracy underscore the need for optimization strategies to improve its feasibility in such contexts.
The final graph compares the S across models, integrating both energy efficiency and predictive performance. Centralized models generally achieve higher sustainability due to their lower energy requirements and superior accuracy. Federated models, however, show varying levels of sustainability depending on the architecture, emphasizing the importance of selecting the appropriate paradigm based on application-specific trade-offs among energy, accuracy, and sustainability.
Additionally, from Table 5.1.3, we observe that in the federated setup, the CNN model achieves the most accurate predictions. Its performance is comparable to that of the neural network architecture with Alpha neurons, which outperform the traditional MLP model.
Furthermore, predictions made with Lapicque neurons are significantly less accurate, with an NRMSE approximately twice as high as that of the CNN model.
The Leaky neuron also performs competitively in the federated setup, achieving a lower NRMSE than the MLP, although its error remains slightly higher than that of the CNN.
Figure 4: Federated Learning Results
[…]
Federated Cumulative results on NRMSE, Energy Consumption and Sustainability of considered models. Model NRMSE Consumption (Wh) $\mathcal{S}$ MLP 3.3134 0.28 29.86 CNN 2.9999 0.64 32.410 Echo State (RC) 4.9471 0.105 36.034 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input 7 spikes/input 10 spikes/input Leaky 3.3061 3.4012 0.0138 0.0213 27.6385 27.9062 Lapicque 6.141 5.131 0.751 0.939 39,1163 38,4670 RLeaky 3.332 3.101 0.832 0.812 33,6647 32.9419 Synaptic 3.621 3.231 0.891 1.234 34,7515 35,6637 Alpha 3.2813 3.1023 1.902 2.523 39.0313 41,0282
5.1.4 Federated vs. Centralized
From Figures 5 and 6, each contain 4 graphs, we observe that in our simulation, federated power consumption is significantly lower compared to centralized configurations. However, this reduction in power consumption comes at the cost of a noticeably higher NRMSE. This trade-off highlights a fundamental challenge in federated learning, where decentralized training introduces communication overheads and reduced data accessibility, leading to diminished predictive accuracy.
Furthermore, for the federated configuration, we find that all models, except the Leaky model, achieve better NRMSE performance with 10 timesteps. In contrast, the Leaky model demonstrates improved metrics with 7 timesteps. This observation underscores the importance of selecting an optimal timestep configuration to balance computational complexity, energy efficiency, and predictive accuracy.
The first graph illustrates the energy consumption of centralized and federated models. Centralized configurations demonstrate consistently lower energy consumption due to localized computations, whereas federated models exhibit higher energy usage owing to the communication overhead inherent in decentralized training.
The second graph compares the predictive performance of the models using the Normalized Root Mean Square Error (NRMSE) metric. Centralized models achieve superior accuracy, reflected in their lower NRMSE values, due to their direct access to the entire dataset. Federated models, while showing higher NRMSE values, maintain reasonable performance levels, making them viable for applications where privacy and decentralization are prioritized over predictive precision.
The third graph visualizes the trade-off between energy consumption and predictive accuracy, with a scatter plot of energy consumption versus NRMSE. Points closer to the origin (0, 0) indicate better resource efficiency and lower predictive error. Centralized models cluster near the origin, showcasing superior energy efficiency and predictive accuracy, while federated models demonstrate a clear trade-off with higher energy costs and slightly reduced accuracy.
The final graph compares the sustainability index (S), which integrates energy efficiency and predictive performance. Centralized models achieve higher sustainability scores due to their lower energy consumption and better accuracy. Federated models exhibit varying levels of sustainability, depending on the architecture and configuration, underscoring the importance of evaluating application-specific trade-offs between energy, accuracy, and sustainability.
Together, these findings emphasize the strengths and limitations of centralized and federated learning approaches. They also highlight the importance of carefully selecting configurations, such as timestep lengths and model architectures, to achieve an optimal balance between energy efficiency, predictive performance, and sustainability in diverse operational contexts.
We observe that energy consumption for federated learning is lower in our simulations than centralized, but the NRMSE values are higher. This indicates that centralized models are more sustainable, as they achieve a lower Sustainability Index (S) compared to federated models. This observation is further supported by the Sustainability Index Comparison Graph, which directly compares the S values of federated and centralized configurations.
In the Energy vs. NRMSE Trade-off graph from the same figures, the blue dots (representing centralized models) exhibit much lower NRMSE values. In this graph, dots closer to the origin (0, 0) represent models with lower S, given the constant dataset size. It is evident from this analysis that the most sustainable models across all experiments are the centralized Leaky models with 7 timesteps and 10 timesteps. Among these, the centralized Leaky model with 7 timesteps is the most sustainable, achieving an S value of 22.960.
Figure 5: Federated vs Centralized Learning 7 timesteps.
Figure 6: Federated vs Centralized Learning 10 timesteps.
6 Discussion and Open Issues
We observe significant variations in power consumption across models during training. For example, the Alpha Neuron architecture for SNNs achieves lower prediction errors but at the cost of substantially higher power consumption. Notably, it demonstrates the best predictive capacity in our centralized training experiments.
These findings highlight that there is no universal rule ensuring bio-inspired models inherently consume less power while maintaining low predictive error. Consequently, a clear trade-off cannot be universally established; each case must instead be analyzed individually.
We do, however, observe a general trend: increased computational complexity tends to result in higher power consumption alongside improved predictive performance. The challenge, therefore, is to select a model with a level of computational complexity that balances acceptable prediction error and manageable power consumption for a specific application.
In our application—cellular traffic forecasting—this consideration is especially critical. The vast amounts of data generated daily raise concerns about significant CO2 emissions resulting from large-scale computations.
At the same time, we must acknowledge the limitations of current models in accurately simulating brain-like learning mechanisms. While the Alpha neuron achieves better predictive capacity by incorporating spiking elements that mimic brain functionality, its internal structure resembles a filter more than the biological structure of a neuron. This raises an intriguing question: how does the human brain achieve such remarkable learning capabilities with such low power consumption?
This question remains unresolved. Meanwhile, the Synaptic neuron, which more closely resembles the functionality of brain cells, exhibits higher error rates compared to classic models but performs better than the Alpha neuron. Additionally, the Synaptic neuron consumes significantly more power than classic models, though less than the Alpha neuron.
The increased power consumption of the Synaptic neuron arises from its attempt to simulate biological processes more accurately, resulting in more complex computations. This suggests there may be a limit to how closely we should model brain functionality. In our case, the more biologically accurate Synaptic neuron produced more erroneous results compared to the Leaky neuron, a simplified, classic Leaky-Integrate-and-Fire neuron.
These observations lead to two important considerations:
  1. 1.
    Specialized versus general-purpose hardware: The cells in our brain are specialized to perform specific functions, whereas our experiments are conducted on general-purpose hardware. This underscores the need to replicate these experiments on neuromorphic hardware designed specifically for each model. Such hardware would allow us to fully exploit their advantages while minimizing disadvantages. Furthermore, specialized hardware would provide more accurate estimations of power consumption, as computational simulations are prone to distortion.
  2. 2.
    Task alignment with brain-like models: Tasks involving numerical pattern recognition and approximate prediction may not align well with the human brain’s natural capabilities. It may be unreasonable to expect a model designed to mimic the human brain to excel in such tasks without question. This does not diminish the learning power of the human brain but instead suggests that we should focus on identifying tasks where these models perform best, reflecting the strengths of human brain functionality.
In addition, the power consumption reported for federated learning in our experiments does not account for the energy required for communication, as these experiments were conducted in a simulated environment. This limitation highlights the need for further studies incorporating real-world network conditions. Such studies should evaluate the true energy footprint of federated learning, particularly in scenarios with varying communication overheads and device heterogeneity.
7 Conclusion
In conclusion, bio-inspired models such as Leaky and Echo State demonstrate significant potential for promoting sustainable machine learning. Their inherent energy efficiency and scalability position them as strong candidates for a wide range of applications, particularly in resource-constrained environments. With further adaptations and optimizations, these models can be refined to achieve superior performance in specific tasks, enhancing their practicality for real-world deployment. Within the centralized and decentralized configuration, our findings highlight the Leaky model as the most sustainable option, offering a compelling solution for energy-efficient machine learning.